OpenStudy (anonymous):

2x + y = 8 x + y = 4 The lines whose equations are given intersect at. A) (4, 0) B) (0, 4) C) all points on the line

OpenStudy (anonymous):

do you know how to use elimination? or substitution? you can use these methods to solve for the intersection of the lines

OpenStudy (anonymous):

@Data_LG2 I'll give ya a medal to help.

OpenStudy (anonymous):

let's eliminate the variable "y" first to solve for "x" by subtracting the first equation to the second equation can you do that for me?

OpenStudy (anonymous):

2x + y = 8 -(x + y = 4) x+0=4 x=4 x + y = 4 4+y=4 y=0 2x + y = 8 [(2)(4)]+0=8 8+0=8 8=8 (A) (4,0)???

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

looks like you're right! (^_^)

OpenStudy (anonymous):

Awesome!!! Thanks, @Data_LG2

OpenStudy (anonymous):

no problem :) keep up the good work! @ihatealgebraplease

OpenStudy (anonymous):

Would you mind helping with this problem @Data_LG2 3x + 2y = 5. 5x + 2y = 7. Based on the given system of equations, which of the following is NOT true. A) 2x = 2 B) 8x = 12 C) 4y = 4

OpenStudy (anonymous):

same thing as you did before, try to solve for x and y if you already found this variables check out which one of the choices is true for a, it says that x is equal to 1 b, x=3/2 c. y=1